$SPAD/src/input free.input

Timothy Daly December 3, 2016

Abstract

We expect that −b < z < a < b

Contents

2

— * —

)set break resume
)spool free.output
)set message test on
)set message auto off
)clear all
--S 1 of 8
Z2:=FreeAbelianGroup Symbol
--R
--R   (1)  FreeAbelianGroup(Symbol)
--R                                                                 Type: Domain
--E 1
--S 2 of 8
a:=a::FreeAbelianGroup Symbol
--R
--R   (2)  a
--R                                               Type: FreeAbelianGroup(Symbol)
--E 2
--S 3 of 8
b:=b::FreeAbelianGroup Symbol
--R
--R   (3)  b
--R                                               Type: FreeAbelianGroup(Symbol)
--E 3
--S 4 of 8
z:=0::FreeAbelianGroup Symbol
--R
--R   (4)  0
--R                                               Type: FreeAbelianGroup(Symbol)
--E 4
--S 5 of 8
a < -b
--R
--R   (5)  false
--R                                                                Type: Boolean
--E 5
--S 6 of 8
-b < z
--R
--R   (6)  true
--R                                                                Type: Boolean
--E 6

--S 7 of 8
z < a
--R
--R   (7)  true
--R                                                                Type: Boolean
--E 7
--S 8 of 8
a < b
--R
--R   (8)  true
--R                                                                Type: Boolean
--E 8
)spool
)lisp (bye)

———-

References^{[1] nothing}

5